# DACoN: DINO for Anime Paint Bucket Colorization with Any Number of Reference Images

### v1

@@ 4.3 Keyframe Colorization
We use the reported results for both BasicPBC and BasicPBC-Ref from the BasicPBC-Ref paper [6]. Since inference code for BasicPBC-Ref is not publicly available, we include only BasicPBC in our qualitative comparison. We conduct our own experiments on the generation-based baselines using their official pre-trained weights. As image and video generation-based methods operate at the pixel level, we apply a post-processing step (Figure 5) to enable segment-level comparison. The detailed colorization process for each method is as follows. For ColorFlow, we use weights optimized for sketches and convert the target image to grayscale before input. For MangaNinja, we disable user-provided correspondence points and use only the grayscaled target image as input. For AniDoc, as it is a video generation method that requires at least 14 conditioning frames, we provide a sequence of 14 binarized target images (all identical) and use the first generated frame for evaluation.
@@ 4.4 Consecutive Frame Colorization
We compare DACoN with previous segment correspondence based approaches designed for consecutive frame colorization [6, 4]. Although the training code for AnT [4] is not publicly available, an official implementation is provided in the Cadmium application [1], which we use for inference. For generation-based methods (i.e., AniDoc [21] and LVCD [10]), we follow the same setup used in keyframe colorization by inputting 14 target images (identical) and using the first generated frame for evaluation.
@@ 4.5 Ablation Study
The quantitative results (see Tables 2 and 4) demonstrate that the DINO-guided Feature Consistency Loss significantly improves accuracy in both keyframe and consecutive frame colorization, with a particularly notable improvement in keyframe colorization. This improvement can be attributed to the ability of DINOv2 to better preserve semantic features when guided by the consistency loss. Furthermore, as illustrated in Figure 9, when color information is removed from the line drawings, shading and highlights in critical areas such as the hair and eyes become less distinct. This highlights the importance of color information for distinguishing segments with similar shapes and positions.
@@ Appendix B Inference Time and Memory Usage
To address this issue, we apply max pooling to downscale the segment masks to a resolution of $512\times 512$, consistent with training conditions, and evaluate the performance under this setting (see “Ours*” in Tables 6 and 7). As a result, we reduce the peak memory usage to approximately 3 GB—about half—without significantly sacrificing colorization accuracy. This indicates that the proposed method can perform robust automatic colorization under memory constraints regardless of input resolution, further enhancing its practicality.

### v2

@@ 4.2 Implementation Details
ColorFlow 1 12.10 13.13 53.75 7.51 58.22
MangaNinja 1 17.39 19.52 8.68 34.51 0.01
AniDoc 1 23.98 27.06 78.71 50.25 86.64
BasicPBC 1 54.30 90.12 66.77 97.28
BasicPBC-Ref 1 63.46 95.31 82.69 98.45
Ours 1 70.61 75.15 97.82 92.16 99.66
ColorFlow 5 13.19 14.83 52.89 11.58 58.88
BasicPBC-Ref 5 64.59 96.12 83.17 98.67
Ours 5 75.49 79.55 98.43 94.72 99.60
Ours max 76.66 80.68 98.60 95.04 99.68
Table 2: Quantitative results of keyframe colorization on the full test set.
Method Ref. shot Acc Acc-Thresh Pix-Acc Pix-F-Acc Pix-B-MIoU
@@ 4.3 Keyframe Colorization
We use the reported results for both BasicPBC and BasicPBC-Ref from the BasicPBC-Ref paper [6]. Since BasicPBC-Ref was evaluated on only 8 of the 10 test characters due to its data requirements, Table 1 shows comparisons on this restricted set, while Table 2 includes results over all test characters. [^1] In the ICCV camera-ready version, only the results of BasicPBC and BasicPBC-Ref were reported based on a subset of the test set. This detail was not explicitly mentioned in the original papers [7, 6], and we found this after checking the released code. For fair comparison, we now provide two tables: Table 1 reports results on the subset, and Table 2 reports results on the full test set (for consistency with the ICCV version). We conduct our own experiments on the generation-based baselines using their official pre-trained weights. As image and video generation-based methods operate at the pixel level, we apply a post-processing step (Figure 5) to enable segment-level comparison. The detailed colorization process for each method is as follows. For ColorFlow, we use weights optimized for sketches and convert the target image to grayscale before input. For MangaNinja, we disable user-provided correspondence points and use only the grayscaled target image as input. For AniDoc, as it is a video generation method that requires at least 14 conditioning frames, we provide a sequence of 14 binarized target images (all identical) and use the first generated frame for evaluation.
@@ 4.4 Consecutive Frame Colorization
We compare DACoN with previous segment correspondence based approaches designed for consecutive frame colorization [6, 4]. Although the training code for AnT [4] is not publicly available, an official implementation is provided in the Cadmium application [1], which we use for inference. Specifically, version 0.3.1 of Cadmium was used for quantitative comparisons, while version 0.3.4 was used for qualitative comparisons. For generation-based methods (i.e., AniDoc [21] and LVCD [10]), we follow the same setup used in keyframe colorization by inputting 14 target images (identical) and using the first generated frame for evaluation.
@@ 4.5 Ablation Study
The quantitative results (see Tables 3 and 5) demonstrate that the DINO-guided Feature Consistency Loss significantly improves accuracy in both keyframe and consecutive frame colorization, with a particularly notable improvement in keyframe colorization. This improvement can be attributed to the ability of DINOv2 to better preserve semantic features when guided by the consistency loss. Furthermore, as illustrated in Figure 9, when color information is removed from the line drawings, shading and highlights in critical areas such as the hair and eyes become less distinct. This highlights the importance of color information for distinguishing segments with similar shapes and positions.
@@ Appendix B Inference Time and Memory Usage
To address this issue, we apply max pooling to downscale the segment masks to a resolution of $512\times 512$, consistent with training conditions, and evaluate the performance under this setting (see “Ours*” in Tables 7 and 8). As a result, we reduce the peak memory usage to approximately 3 GB—about half—without significantly sacrificing colorization accuracy. This indicates that the proposed method can perform robust automatic colorization under memory constraints regardless of input resolution, further enhancing its practicality.